OpenStudy (anonymous):

@mathemathics In squareroot x+4=5 @mathemathics In squareroot x+4=5 @Mathematics

OpenStudy (anonymous):

\[ln (\sqrt{x+4})=5\] \[e^{5}=\sqrt{x+4}\] \[(e^5)^2=x+4\] \[e^{10}=x+4\] \[x=e^{10}-4\]

OpenStudy (anonymous):

yes that is it